OpenStudy (anonymous):

A package contains 6 candy canes, 4 of which are cracked. If 2 are selected, find the probability of getting exactly two cracked candy canes.

OpenStudy (anonymous):

see total ways of selecting two candies are C(6,2) =15 now no. of ways of fav outcome is C(4,2) = 6 thus reqd prob=6/15 =2/5

OpenStudy (anonymous):

Thanks

OpenStudy (anonymous):

welcome dear